# Neural Posterior Estimation for Cataloging Astronomical Images with Spatially Varying Backgrounds and Point Spread Functions

### v1

@@ 6 Design of Numerical Experiments
To generate synthetic datasets for training and validation, we sampled two variants of our prior distribution. Both variants adhere to the description in Section 4.1 but differ in the collection of SDSS fields they draw on to define the empirical distributions of backgrounds and PSFs. One prior, called $p_{\text{one}}$, samples the background and PSF from a single SDSS field (run 94, camcol 1, field 12); it serves as a baseline. The other prior, called $p_{\text{many}}$, samples from $500$ SDSS fields. As can be seen in Figure 2, the field $p_{\text{one}}$ is based on is fairly typical in terms of PSF width or background level: $p_{\text{many}}$ has support for fields with wider and narrower PSFs and brighter and fainter background levels.
@@
Single-field from a single SDSS field not provided
PSF-unaware from many SDSS fields not provided
PSF-aware from many SDSS fields encoding provided
@@
Each network is trained with 210k simulated images for 50 epochs using the Adam optimizer and an initial learning rate of $1\times 10^{-3}$, which decays by a factor of 0.1 after 32 epochs. Early stopping regularization is based on a 30k image held-out validation image set. Training times are similar for these three networks, with each network requiring roughly 10 hours to train using a single NVIDIA 2080 Ti GPU. We compare our models using various metrics, including precision and recall in detection, star/galaxy separation accuracy, and flux measurement error.
Ideally, an inference network is robust to modest differences in the training and testing data distributions. However, deep neural networks often perform poorly if the training and testing data distributions differ (Zhou et al., 2022). The canonically “easiest” scenario for a network is to evaluate images drawn from the distribution on which the network has trained. We therefore expect that the single-field network will perform well when evaluated on images sampled from $p_{\text{one}}$ and poorly with images sampled from $p_{\text{many}}$.
In our experiments, we quantify the relative improvements to network robustness in the case where the network first learns the distribution across many SDSS fields and in the further case in which the network explicitly makes use of the image-specific background and PSF. For simplicity, we restrict our analysis to the r band in SDSS, which corresponds to a wavelength centered at 6166 angstroms (Stoughton et al., 2002).
In plots of our experimental results, we generally stratify/bin performance by objects’ true r-band magnitudes. For performance stratified by signal-to-noise ratio (SNR) rather than magnitude, see Appendix A. The widths of the bins are selected so that each bin contains an equal number of objects. The shaded regions indicate the 1-$\sigma$ deviation according to bootstrap resampling of $80\times 80$-pixel images in a test set containing 42k images (and corresponding ground-truth catalogs). The width of the shaded region indicates how the performance of a particular network may differ if the experiments were repeated with new testing data, but it is too conservative to indicate whether the relative performance difference is statistically significant, as all methods are tested with the same test set (a “matched pairs” experimental design). The bracketed values indicate the average performance across all bins for each network.
7 Results
@@ 7.1 Diverse Backgrounds and PSFs
We first compare our three inference networks using synthetic data sampled from the $p_{\text{many}}$ prior. To match centroids of detected sources to sources in the actual catalog, we use the Jonker-Volgenant algorithm to find a bipartite matching that minimizes the sum of the distances between true and predicted light sources (Jonker & Volgenant, 1987).
@@ 7.1.1 Light Source Detection
We first compare the performance of our inference networks for light source detection. Figures 4(a) and 4(b) show the precision (purity) and the recall (completeness) of the detections. The precision is binned by the estimated source r-band magnitude, while the recall (as with all other plots in this section) is binned by the true source r-band magnitude. As expected, the detection performance of all three networks increases monotonically for brighter sources. At almost all magnitudes, the PSF-unaware and PSF-aware networks, which are shown with red squares and orange triangles, respectively, outperform the single-field network, which is shown with blue circles. The exception is in the detection precision for sources in the second-faintest bin, where the networks perform comparably well (within the 1-$\sigma$ bootstrap interval). Interestingly, the addition of PSF information to evaluation images through parameter-based encoding in the PSF-aware network does little to improve source detection at any magnitude. This implies that to mitigate false detections and detect sources fainter than an r-band magnitude 21.23, the network most benefits from seeing a wider distribution of backgrounds and PSFs; it does not need access to the image-specific background and PSF for robust performance in source detection.
@@
Figure 4: Performance for held-out data with a varied backgrounds and PSFs. Solid lines indicate the mean performance in each r-band magnitude bin. Bin boundaries are chosen so that each bin has an equal number of evaluation points. Shaded regions indicate the 1-$\sigma$ deviation according to the bootstrap. The bracketed values indicate the average performance across all bins for each inference network.
@@ 7.1.2 Star/Galaxy Separation
In Figure 4(c), we show the accuracy of the star/galaxy separation of light sources. If the posterior source-type probability (see Table 2) for an object is greater than 0.5, we classify it as a galaxy; otherwise, we classify it as a star. We further investigate the choice of this classification threshold in Section 7.3.2. Again, at all r-band magnitudes shown, the PSF-unaware and PSF-aware networks outperform the single-field network in accuracy.
For this task, the PSF-aware network outperforms the PSF-unaware network for r-band magnitudes greater than 21, but has comparable performance at other magnitudes. Note that most objects with r-band magnitudes $<19$ are likely to be stars, whereas those with r-band magnitudes $>21$ are likely to be galaxies. This suggests that the explicit incorporation of PSF information may aid in identifying galaxies more than stars. Interestingly, the single-field network does not exhibit monotonic behavior with source magnitude. Instead, its classification performance decreases at intermediate magnitudes 19 and 21.5, where we might expect star/galaxy mixing in a magnitude range where the classes are relatively more balanced than at the extremes. The PSF-unaware and PSF-aware networks mitigate this mixing.
@@ 7.1.3 Flux Measurement
Figure 4(d) compares performance for flux prediction, quantified by the mean absolute percentage error (MAPE):
@@
We find that both the PSF-unaware and PSF-aware networks outperform the single-field network at all magnitudes. The PSF-unaware network is marginally better than the PSF-aware network for the brightest objects (r-band magnitude $<17.78$), and the PSF-aware network is marginally better for fainter objects (r-band magnitude 21.23–21.75).
All three models are generally better at estimating the flux of brighter objects, with increasing error at higher magnitudes. The one exception to this is the single-field model in the brightest bin (r-band magnitude $<17.78$). As most objects in this magnitude bin tend to be stars (which look like the PSF), this suggests that seeing a variety of PSFs during training aids in improving flux measurement of these sources in particular.
@@ 7.2 Similar Backgrounds and PSFs
Figure 5 shows the same metrics for detection, star/galaxy separation, and flux measurement as in Section 7.1, but now evaluated with this new data distribution. In all tasks, the PSF-unaware and PSF-aware networks are competitive with the single-field network, despite being evaluated on an out-of-distribution test dataset. In fact, all three models are nearly identical at low ($<21$) magnitudes for all three metrics. The PSF-unaware network has slightly poorer precision, classification accuracy, and flux MAPE at fainter magnitudes than the PSF-aware and single-field models, which are consistently within 1-$\sigma$ of each other.
@@
Figure 5: Performance for held-out data with a fixed background and PSF. Solid lines indicate the mean performance in each r-band magnitude bin. Bin boundaries are chosen so that each bin has an equal number of evaluation points. Shaded regions indicate the 1-$\sigma$ deviation according to the bootstrap. The bracketed values indicate the average performance across all bins for each inference network. In contrast to the setting of Figure 4, here all networks perform similarly.
@@ 7.3.1 Detection
Figure 6: The expectation number of detections vs. the true number of sources. Left: all magnitudes. Center: the brightest sources ($<$17.78 magnitude). Right: the faintest sources (21.75–22 magnitude). Each point represents the average over all images with a particular number of sources. The diagonal dotted gray line represents perfect performance, where the number of sources always equals the number of detections. All three networks perform equally well on bright sources and are underconfident on faint sources.
Figure 6 shows calibration for source detection. The figure panels show the expected number of sources in an 80 $\times$ 80 pixel image, given by summing the probabilities of detections, as a function of the actual number of sources in that image. From left to right, we show this for all sources, the top 10th percentile of the brightest sources ($<17.78$ magnitude), and the bottom 10th percentile of the faintest sources (21.7521.7521.7521.7522222222 magnitude). The dashed gray line traces the behavior of a perfectly calibrated network, which would detect the exact number of sources.
Overall, all three networks are fairly well-calibrated; that is, the expected value of detected sources is almost equal to the number of actual sources, with some underconfidence when there are many sources in an image ($>8$).
@@
When restricting the analysis to the faintest sources in our dataset (with r-band magnitude between 21.75 and 22), all three networks are underconfident in their predictions, systematically predicting fewer faint objects than truth. This is also consistent with the intuition that faint objects are the most difficult to detect. There is a minimal difference between the calibration of the PSF-unaware and PSF-aware networks. This is consistent with the detection metrics in Figures 4(a) and 4(b), where explicit PSF information also led to limited improvement in overall detection performance.
@@ 7.3.2 Star/Galaxy Separation
Figure 7: Receiver operating characteristic (ROC) curve for star/galaxy separation. The true positive rate for galaxies (sensitivity) is plotted against the false positive rate (1 - specificity) for different classification thresholds. Left: all magnitudes. Center: the brightest sources (magnitude $<$17.78). Right: the faintest sources (magnitude 21.75–22). The dotted gray line represents the performance of a random classifier. The area under the curve (AUC) for each inference network is provided in brackets.
@@
We find that the PSF-unaware and PSF-aware models significantly outperform the single-field network at all magnitudes, although the AUC for all three models is respectable. At bright magnitudes, the PSF-unaware and PSF-aware models achieve near-perfect discrimination, consistent with Figure 4(c). At faint magnitudes, the difference between models is more pronounced, with the PSF-aware network showing a slight improvement over the PSF-unaware network.
@@ 7.3.3 Flux Measurement
Figure 8: Proportion of sources whose flux falls within various credible intervals. Left: all magnitudes. Center: the brightest sources ($<$17.78 magnitude). Right: the faintest sources (21.75–22 magnitude). The dotted gray line represents perfect performance, where the credible interval contains exactly that percentage of sources.
In this section, we examine how well calibrated our networks are when estimating the flux of each source by comparing the fraction of true flux values contained in the credible interval of the posterior output by our networks. Credible intervals are the Bayesian analogue of confidence intervals; they describe the probability that the true flux should fall within that interval. For example, for a perfectly calibrated network, a 95% credible interval contains the true flux 95% of the time.
@@
The single-field network is underdispersed for all magnitude ranges, meaning that the variance predicted by the network is lower than the observed variance. An underdispersed network produces overly narrow credible intervals, which would contain the true flux less often than expected. For example, when evaluated on all sources, the 60% credible intervals of the single-field network contain only 48% of the true flux values.
@@
We find that the PSF-aware network is more poorly calibrated than the PSF-unaware network for all sources, but better calibrated for the brightest and faintest sources. Furthermore, all three networks are best calibrated for faint sources and more poorly calibrated for bright sources. This suggests that although predicting the center of the distribution is easier for the brightest objects (see Figure 4(d)), uncertainty estimation is more error-prone at these magnitudes.
This miscalibration is likely due to the restrictive parametric form of the variational distribution, which approximates flux as log normal. We conjecture that for faint light sources the posterior may be relatively diffuse and, therefore, adequately approximated by a log-normal distribution. In contrast, for bright light sources, the data may exert much more influence on the posterior than on the prior, resulting in a posterior that is sharply peaked and fairly symmetric and therefore not well approximated by a log-normal, which has a heavy right tail. A more flexible family of variational distributions, such as a mixture of Gaussians, could lead to better calibration.
@@ 8 Discussion
Training inference networks on simulated images with varied backgrounds and PSFs aids in detection performance by increasing robustness across fields. However, little improvement in detection performance is gained by providing explicit information about the background and PSF to the inference network (see Figures 4(a) and 4(b)).
Providing an inference network with PSF information (e.g., through dedicated input channels) can improve performance in star/galaxy separation and flux measurement. The most noticeable improvements were in star/galaxy separation for objects of intermediate brightness and in flux measurement for the brightest objects (see Figures 4(a), 4(b), and 8).
In all tasks, inference networks trained with a variety of PSFs and backgrounds performed no worse than an inference network trained with a single background and PSF, even when evaluated on images generated with this specific background and PSF (see Figure 5).
Training inference networks with varied PSFs and backgrounds improves calibration for star/galaxy separation and flux measurement. In addition, providing an inference network with explicit information about varied PSFs improves flux calibration, especially for the brightest and faintest sources (see Figure 8).
With PSF information, our results for light source detection and star/galaxy separation are comparable to those reported for other image analysis methods such as Shi et al. (2022); Sevilla-Noarbe et al. (2018); Clarke et al. (2020). Specifically, our detection precision exceeds 0.95 for r-band magnitudes brighter than 21.5, and our detection recall exceeds 0.9 for r-band magnitudes brighter than 21.2. Our star/galaxy separation accuracy exceeds 0.96 for r-band magnitudes brighter than 22.
We were surprised that the single-field network performed as well as it did in our experiments, particularly for detection, given that it was not trained with a variety of PSFs and background intensities (see Figures 4(a), 4(b), and 6). One possible explanation is that the variation in background intensity matters more for detection than the variation in PSF and that the background intensity varies relatively little between fields, as indicated in Figure 2. For tasks that require more knowledge of the shape of objects, such as star/galaxy separation and flux measurement, we observed that incorporation of PSF variation produces more significant improvements.
@@
Enhanced PSF encoding. In this work, we used the six-parameter PSF encoding provided by the SDSS pipeline, which represents the PSF as spherical. However, the true PSF is unlikely to be exactly spherical or otherwise perfectly characterized by six parameters. While the six-parameter PSF encoding may be adequate in some applications, NPE can be particularly vulnerable to model misspecification (Ward et al., 2022). A more sophisticated PSF encoding could be obtained using an auxiliary autoencoder, as in Jia et al. (2020). By allowing an autoencoder to learn a compressed representation of the PSF, we allow for greater expressivity than a simple parametric model.
Implications for deeper surveys. We conducted our experiments with synthetic data designed to mirror SDSS data. Mirroring the data of a well-understood survey simplified the generation of realistic synthetic data. For instance, we could fit our prior distribution to existing catalogs without needing to extrapolate to what sources would be detectable in images of greater depth.
@@
Deblending performance. In Sections 7.1.1 and 7.3.1, we examine the extent to which our method infers the correct number of light sources. This is a particularly challenging problem in crowded fields where light sources overlap and blend. In ongoing work, we are further investigating how detection performance varies with blendedness, in relation to established deblending methods such as Source Extractor (Bertin & Arnouts, 1996) and Scarlet (Melchior et al., 2018). Because this work focuses on spatially varying covariates (background and PSF), we do not explicitly consider the degree of blendedness in our metrics for detection, or for other tasks.
Galaxy shape measurement. To assess the success of our integration of spatially varying covariates (background and PSF) with NPE, we considered three tasks: detection, star/galaxy separation, and flux measurement. We are intrigued by the possibility of using NPE for galaxy-shape measurement in the context of properly modeling PSF. A proper accounting of the PSF should be particularly helpful in estimating the half-light radius. However, we leave this to follow-up work, as our current three tasks seem adequate to demonstrate NPE with spatially variable covariates. One challenge in benchmarking galaxy shape measurement is in defining scientifically relevant performance metrics. If the ultimate use of these galaxy shape measurements is to estimate shear-shear correlation for weak lensing analyses, as is often the case, then average ellipticity error may not be a meaningful measure of error.

### v2

@@ 6 Numerical Experiments
To generate synthetic datasets for training and validation, we sampled two variants of our prior distribution. Both variants adhere to the description in Section 4.1 but differ in the collection of SDSS fields they draw on to define the empirical distributions of backgrounds and PSFs. One prior, called $p_{\text{one}}$, samples the background and PSF from a single SDSS field (run 94, camcol 1, field 12); it serves as a baseline. The other prior, called $p_{\text{many}}$, samples from $500$ SDSS fields. As can be seen in Figure 2, the field $p_{\text{one}}$ is based on is fairly typical in terms of PSF width or background level: $p_{\text{many}}$ has support for fields with wider and narrower PSFs and brighter and fainter background levels. For simplicity, we restrict our data to the r band in SDSS, which corresponds to a wavelength centered at 6166 angstroms (Stoughton et al., 2002).
@@
Single-field a single SDSS field not provided
PSF-unaware many SDSS fields not provided
PSF-aware many SDSS fields encoding provided
@@
Each network is trained with 210k simulated images for 50 epochs using the Adam optimizer and an initial learning rate of $1\times 10^{-3}$, which decays by a factor of 0.1 after 32 epochs. Early stopping regularization is based on a 30k image held-out validation image set. Training times are similar for these three networks, with each network requiring roughly 10 hours to train using a single NVIDIA 2080 Ti GPU.
Our primary question is whether either of the generalist networks (PSF-unaware and PSF-aware) can keep up with the specialist network (single-field) in the setting in which the specialist network specializes: with held-out data sampled from the same population as its training data was. A secondary question is whether providing a generalist network with explicit PSF information as input is necessary: it seems like it would be, but it is possible that PSF information could be learned implicitly from the input image, or that downstream metrics are insensitive to PSF variation. A final question is whether training networks with a variety of PSFs and backgrounds is even necessary, or whether a network trained with a particular PSF and background can perform well out-of-distribution and beyond the support of its training data (that is, on PSFs and backgrounds unlike those it encountered during training). We consider this third question be worthy of verification, but of secondary importance because a priori there is strong evidence that its answer in negatory: neural networks are widely known to perform poorly in the presence of domain shift (Zhou et al., 2022).
@@ 6.1 Diverse Backgrounds and PSFs
We first compare our three inference networks on synthetic data sampled from the $p_{\text{many}}$ prior. We match centroids of detected sources to ground-truth sources by using the Hungarian algorithm (Kuhn, 1955) to find a bipartite matching that minimizes the sum of the distances.222 We do not take a position on whether finding an optimal bipartite match is important. Because the Hungarian algorithm runs efficiently in our setting and is readily available via the scipy Python package, we do not explore inexact alternatives. To ensure fair comparisons, the same matching algorithm is used for all evaluations.
@@ 6.1.1 Light Source Detection
We first compare the performance of our inference networks for light source detection. Figures 4 and 4 show the precision (purity) and the recall (completeness) of the detections. The precision is binned by the estimated source r-band magnitude, while the recall (as with all other plots in this section) is binned by the true source r-band magnitude. As expected, the detection performance of all three networks increases monotonically for brighter sources. At almost all magnitudes, the PSF-unaware and PSF-aware networks, which are shown with red squares and orange triangles, respectively, outperform the single-field network, which is shown with blue circles. The exception is in the detection precision for sources in the second-faintest bin, where the networks perform comparably well (within the 1-$\sigma$ bootstrap interval). Interestingly, the addition of PSF information to evaluation images through parameter-based encoding in the PSF-aware network does little to improve source detection at any magnitude. This implies that to mitigate false detections and detect sources fainter than an r-band magnitude 21.23, the network most benefits from seeing a wider distribution of backgrounds and PSFs; it does not need access to the image-specific background and PSF for robust performance in source detection.
We note that precision and recall remain slightly imperfect even for the brightest bins due to a combination of factors: (1) close pairs of bright sources can blend and be merged or misassigned by the network; (2) our current centroid‐matching uses a strict 1 pixel radius (increasing this to 2–3 pixels would recover additional matches); and (3) as described in Section 5.1, the 4 $\times$ 4 pixel tiling scheme permits at most one detection per tile, so sources straddling tile boundaries are occasionally dropped. We plan to address these limitations in future work.
@@
Figure 4: Performance for held-out data with varied backgrounds and PSFs. Solid lines indicate the mean performance in each r-band magnitude bin. The widths of the bins are selected so that each bin contains an equal number of objects. The shaded regions indicate the 1-$\sigma$ deviation according to bootstrap resampling of images (and corresponding ground-truth catalogs) in a test set containing 42k images, each $80\times 80$ pixels. The bracketed values indicate the average performance of an inference network across all bins. For similar plots with performance stratified by signal-to-noise ratio (SNR) rather than magnitude, see Appendix A.
@@ 6.1.2 Star/Galaxy Separation
In Figure 4, we show the accuracy of the star/galaxy separation of light sources. If the posterior source-type probability (see Table 2) for an object is greater than 0.5, we classify it as a galaxy; otherwise, we classify it as a star. We further investigate the choice of this classification threshold in Section 6.3.2. Again, at all r-band magnitudes shown, the PSF-unaware and PSF-aware networks outperform the single-field network in accuracy.
For this task, the PSF-aware network outperforms the PSF-unaware network for r-band magnitudes fainter than 21, but has comparable performance at other magnitudes. Note that most objects with r-band magnitudes brighter than $19$ are likely to be stars, whereas those with r-band magnitudes fainter than $21$ are likely to be galaxies. This suggests that the explicit incorporation of PSF information may aid in identifying galaxies more than stars. Interestingly, the single-field network does not exhibit monotonic behavior with source magnitude. Instead, its classification performance decreases at intermediate magnitudes 19 and 21.5, where we might expect star/galaxy mixing in a magnitude range where the classes are relatively more balanced than at the extremes. The PSF-unaware and PSF-aware networks mitigate this mixing.
@@ 6.1.3 Flux Measurement
Figure 4 compares performance for flux prediction, quantified by the mean absolute percentage error (MAPE):
@@
We additionally overlay a theoretical lower bound on the flux error, derived from the expected flux uncertainty for isolated sources using real SDSS parameters (gain, background, PSF width, and flux calibration) for run 94, camcol 1, field 12. This bound approximates the minimal achievable MAPE due to photon and background noise.
We find that both the PSF-unaware and PSF-aware networks outperform the single-field network at all magnitudes. The PSF-unaware network is marginally better than the PSF-aware network for the brightest objects (r-band magnitude <17.78<17.78), and the PSF-aware network is marginally better for fainter objects (r-band magnitude 21.23–21.75).
All three models are generally better at estimating the flux of brighter objects, with greater error at fainter magnitudes. The one exception to this is the single-field model in the brightest bin (r-band magnitude <17.78<17.78). As most objects in this magnitude bin tend to be stars (which look like the PSF), this suggests that seeing a variety of PSFs during training aids in improving flux measurement of these sources in particular.
@@ 6.2 Similar Backgrounds and PSFs
Figure 5 shows the same metrics for detection, star/galaxy separation, and flux measurement as in Section 6.1, but now evaluated with this new data distribution. In all tasks, the PSF-unaware and PSF-aware networks are competitive with the single-field network, despite being evaluated on a test dataset sampled from a different distribution than the training set. In fact, all three models are nearly identical at bright (<21<21) magnitudes for all three metrics. The PSF-unaware network has slightly poorer precision, classification accuracy, and flux MAPE at fainter magnitudes than the PSF-aware and single-field models, which are consistently within 1-$\sigma$ of each other.
@@
Figure 5: Performance for held-out data with a fixed background and PSF. As in Figure 4, solid lines indicate the mean performance in each r-band magnitude bin, the widths of the bins are selected so that each bin contains an equal number of objects, the shaded regions indicate the 1-$\sigma$ deviation according to bootstrap resampling, and the bracketed values indicate the average performance of an inference network across all bins. In contrast to the setting of Figure 4, here all networks perform similarly.
@@ 6.3.1 Detection
Figure 6: The expectation number of detections vs. the true number of sources. Left: all magnitudes. Center: the brightest sources (<<17.78 magnitude). Right: the faintest sources (21.75–22 magnitude). Each point represents the average over all images with a particular number of sources. The diagonal dotted gray line represents perfect performance, where the number of sources always equals the number of detections. All three networks perform equally well on bright sources and are underconfident on faint sources.
Figure 6 shows calibration for source detection. The figure panels show the expected number of sources in an 80 $\times$ 80 pixel image, given by summing the probabilities of detections, as a function of the actual number of sources in that image. From left to right, we show this for all sources, the top 10th percentile of the brightest sources (<17.78<17.78 magnitude), and the bottom 10th percentile of the faintest sources (21.7521.752222 magnitude). The dashed gray line traces the behavior of a perfectly calibrated network, which would detect the exact number of sources.
Overall, all three networks are fairly well-calibrated; that is, the expected value of detected sources is almost equal to the number of actual sources, with some underconfidence when there are many sources in an image (>8>8).
@@
When restricting the analysis to the faintest sources in our dataset (with r-band magnitude between 21.75 and 22), all three networks are underconfident in their predictions, systematically predicting fewer faint objects than truth. This is also consistent with the intuition that faint objects are the most difficult to detect. There is a minimal difference between the calibration of the PSF-unaware and PSF-aware networks. This is consistent with the detection metrics in Figures 4 and 4, where explicit PSF information also led to limited improvement in overall detection performance.
@@ 6.3.2 Star/Galaxy Separation
Figure 7: Receiver operating characteristic (ROC) curve for star/galaxy separation. The true positive rate for galaxies (sensitivity) is plotted against the false positive rate (1 - specificity) for different classification thresholds. Left: all magnitudes. Center: the brightest sources (magnitude <<17.78). Right: the faintest sources (magnitude 21.75–22). The dotted gray line represents the performance of a random classifier. The area under the curve (AUC) for each inference network is provided in brackets.
@@
We find that the PSF-unaware and PSF-aware models significantly outperform the single-field network at all magnitudes, although the AUC for all three models is respectable. At bright magnitudes, the PSF-unaware and PSF-aware models achieve near-perfect discrimination, consistent with Figure 4. At faint magnitudes, the difference between models is more pronounced, with the PSF-aware network showing a slight improvement over the PSF-unaware network.
@@ 6.3.3 Flux Measurement
Figure 8: Proportion of sources whose flux falls within various credible intervals. Left: all magnitudes. Center: the brightest sources (<<17.78 magnitude). Right: the faintest sources (21.75–22 magnitude). The dotted gray line represents perfect performance, where the credible interval contains exactly that percentage of sources.
We now examine how well calibrated our networks are when estimating the flux of each source by comparing the fraction of true flux values contained in the credible interval of the posterior output by our networks. Credible intervals are the Bayesian analogue of confidence intervals; they describe the probability that the true flux should fall within that interval. For example, for a perfectly calibrated network, a 95% credible interval contains the true flux 95% of the time.
@@
The single-field network is underdispersed for all magnitude ranges, meaning that the variance predicted by the network is lower than the observed variance. An underdispersed network produces overly narrow credible intervals, which would contain the true flux less often than expected. For example, when evaluated on all sources, the 60% credible intervals of the single-field network contain only 48% of the true flux values. The generalist networks (PSF-aware and PSF-unaware) are somewhat better calibrated than the specialist network (single-field), but the degree to which the former are better is less than the degree to which all networks are miscalibrated. This shared misspecification can happen for a variety of reasons, but in variational inference, the leading reason is restrictions on the form of the variational distribution, which in this case approximates the flux posterior as log normal. This source of miscalibration may be mitigated by a more flexible variational distribution, such as a mixture of normal distribution, but actually doing so is beyond the scope of this work, which is focused on comparing inference network training regimes.
@@
We find that the PSF-aware network is more poorly calibrated than the PSF-unaware network for all sources, but better calibrated for the brightest and faintest sources. Furthermore, all three networks are best calibrated for faint sources and more poorly calibrated for bright sources. This suggests that although predicting the center of the distribution is easier for the brightest objects (see Figure 4), uncertainty estimation is more error-prone at these magnitudes.
@@ 7 Discussion
Training inference networks on simulated images with varied backgrounds and PSFs aids in detection performance by increasing robustness across fields. Training with varied PSFs and backgrounds also improves calibration for star/galaxy separation and flux measurement. In all tasks, inference networks trained with a variety of PSFs and backgrounds performed no worse than an inference network trained with a single background and PSF, even when evaluated on images generated with this specific background and PSF (see Figure 5).
Providing an inference network with PSF information (e.g., through dedicated input channels) can improve performance in star/galaxy separation and flux measurement. The most noticeable improvements were in star/galaxy separation for objects of intermediate brightness and in flux measurement for the brightest objects (see Figures 4, 4, and 8). These results for light source detection and star/galaxy separation, obtained with explicit PSF information, are comparable to those reported for other image analysis methods such as Shi et al. (2022); Sevilla-Noarbe et al. (2018); Clarke et al. (2020). Specifically, our detection precision exceeds 0.95 for r-band magnitudes brighter than 21.5, and our detection recall exceeds 0.9 for r-band magnitudes brighter than 21.2. Our star/galaxy separation accuracy exceeds 0.96 for r-band magnitudes brighter than 22. However, little improvement in detection performance is gained by providing explicit information about the background and PSF to the inference network (see Figures 4 and 4).
Although the single-field network was often outperformed by the other networks, we were surprised that it performed as well as it did in our experiments, particularly for detection, given that it was not trained with a variety of PSFs and background intensities (see Figures 4, 4, and 6). The conventional wisdom is that neural networks are extremely brittle when faced with inputs outside of the support of the distribution of their training data, and indeed, we observed some evidence of this in that the single-field network was always outperformed in these settings. However, this conventional wisdom lacks nuance about how far beyond the support of the training distribution a particular input is. For our particular network architecture and the degree of PSF and background variation in observed in SDSS, this conventional wisdom, though broadly correct, may overstate the concern.
@@
Enhanced PSF encoding. In this work, we used the six-parameter PSF encoding provided by the SDSS pipeline, which represents the PSF as isotropic. However, the true PSF is unlikely to be exactly isotropic or otherwise perfectly characterized by six parameters. While the six-parameter PSF encoding may be adequate in some applications, a more sophisticated PSF encoding could be obtained using an auxiliary autoencoder, as in Jia et al. (2020). By allowing an autoencoder to learn a compressed representation of the PSF, we allow for greater expressivity than a simple parametric model.
Application to deeper surveys. We conducted our experiments with synthetic data designed to mirror SDSS data. Mirroring the data of a well-understood survey simplified the generation of realistic synthetic data. For instance, we could fit our prior distribution to existing catalogs without needing to extrapolate to what sources would be detectable in images of greater depth.
@@
Quantifying deblending performance. In Sections 6.1.1 and 6.3.1, we examine the extent to which our method infers the correct number of light sources. This is a particularly challenging problem in crowded fields where light sources overlap and blend. In ongoing work, we are further investigating how detection performance varies with blendedness, in relation to established deblending methods such as Source Extractor (Bertin & Arnouts, 1996) and Scarlet (Melchior et al., 2018). Because this work focuses on spatially varying covariates (background and PSF), we do not explicitly consider the degree of blendedness in our metrics for detection or for other tasks.
Mitigating model misspecification. Imperfect simulators pose two distinct problems for neural posterior estimation (NPE). First, like most inference methods, simulation mismatch yields a posterior that is inaccurate proportional to the misspecification. Second—and specific to NPE—the inference network may be presented at inference time with out‑of‑distribution data.
In this work, we sidestep both issues by training and evaluating exclusively on synthetic data from the same simulator, to address the challenges of spatially varying covariates in isolation of issues of model misspecification. In the application of NPE to real astronomical images, some degree of model misspecification is inevitable. Previous work applying NPE to crowded starfields, using real SDSS images, did not require any special techniques to overcome model misspecification beyond carefully tuning the simulator (Liu et al., 2023b). However, a variety of techniques have been developed to mitigate model misspecification in NPE. For example, Ward et al. (2022) shows that naïve NPE can produce unreliable posterior approximations under misspecification and proposes to model a simulator–target error channel that “denoises” observations and recovers credible posterior estimates. Huang et al. (2023) shows that summary‑statistic regularization in NPE suppresses simulator‑sensitive features and stabilizes both embeddings and posteriors under mismatch.
Finally, we note that inputting imperfect pipeline estimate of PSF into our PSF-aware inference networks is largely a separate issue from model misspecification, and one that poses little challenge for our methods. During training, an inference network naturally learns to adjust or ignore misleading inputs, without requiring any explicit domain adaptation.
Measuring galaxy shapes. To assess the success of our integration of spatially varying covariates (background and PSF) with NPE, we considered three tasks: detection, star/galaxy separation, and flux measurement. We are intrigued by the possibility of using NPE for galaxy-shape measurement in the context of properly modeling PSF. A proper accounting of the PSF should be particularly helpful in estimating the half-light radius. However, we leave this to follow-up work, as our current three tasks seem adequate to demonstrate NPE with spatially variable covariates. One challenge in benchmarking galaxy shape measurement is in defining scientifically relevant performance metrics. If the ultimate use of these galaxy shape measurements is to estimate shear-shear correlation for weak lensing analyses, as is often the case, then average ellipticity error may not be a meaningful measure of error.